find the smallest sphere which touches the lines (x-2)/1=(y-1)/-2=(z-6)/1 and (x+3)/7=(y+3)/-6=(z+3)/1 


1
Expert's answer
2020-09-20T17:51:37-0400
"Solution"

We will first find the distance between the two lines. Then that is the diameter of smallest sphere touching the two st. lines. Dividing the diameter by 2 we will get the radius of the sphere.


The two lines "L_1" and "L_2" are

"<t_1-2,\ 2-2t_1,\ t_1-6>\ and\ <7t_2+21,\ -6t_2-18,\ t_2+3>"

The vector "\vec{P}" between the two touchpoints "P _1" and "P _2" must be perpendicular to both direction vectors of the lines "L_1\ and\ L_2"


"\vec{P}=P_1-P_2=
\begin{bmatrix}
(t_1-2)-(7t_2+21)\\
(-2t_1+2)-(-6t_2-18)\\
(t_1-6)-(t_2+3)
\end{bmatrix} \implies 
\begin{bmatrix}
t_1-7t_2-23\\
-2t_1+6t_2+20\\
t_1-t_2-9
\end{bmatrix}\\

\vec{P}\begin{bmatrix}1\\ -2\\ 1 \end{bmatrix}=(t_1-7t_2-23)-2(-2t_1+6t_2+20)+(t_1-t_2-9)=0\\
\implies 6t_1-20t_2-72=0\\
\vec{P}\begin{bmatrix}7\\ -6\\ 1 \end{bmatrix}=-7(t_1-7t_2-23)+6(-2t_1+6t_2+20)+(t_1-t_2-9)=0\\
\implies -20t_1+86t_2+290=0\\"

Solving simultaneously, we get

"t_1=\frac{98}{29}\\
t_2=-\frac{75}{29}\\
P_1=\{\frac{40}{29}\, -\frac{138}{29}, \ -\frac{76}{29} \}\\
P_2=\{\frac{84}{29}\, -\frac{72}{29}, \ \frac{12}{29} \}\\"

The center of the sphere is the middle between the touchpoints:


"M=\{\frac{62}{29},\ -\frac{105}{29},\ -\frac{32}{29},\ \}\\
r^2=\frac{|\vec{P}|}{4}=\frac{(t_1-7t_2-23)^2+(-2t_1+6t_2+20)^2+(t_1-t_2-9)^2}{4}=\frac{121}{29}"


Thus the smallest sphere is the sphere with the equation


"(x-\frac{62}{29})^2+(y+\frac{105}{29})^2+(z+\frac{32}{29})^2=\frac{121}{29}--->answer"